OpenStudy (anonymous):

how do you figure out the area of a circle with only a diameter of 1 mile?

OpenStudy (4n1m0s1ty):

area of circle = pi*r^2 where r is diameter/2

OpenStudy (anonymous):

are = pi*r^2 r = radius = diameter / 2